factor completely

  1. 400-25y2
  2. x3+x2+x+1
  3. x6-7x3-8
  4. 16m2-56mn+49n2
  5. (x2+1)-1/2(x-7)2(x2+1)1/2(x-7)
1
Expert's answer
2021-03-16T07:11:40-0400

1.


"400-25y^2=25(16-y^2)=25(4-y)(4+y)"

2.


"x^3+x^2+x+1=x^2(x+1)+(x+1)"

"=(x+1)(x^2+1)"

3.


"x^6-7x^3-8=(x^3+1)(x^3-8)"

"=(x+1)(x-2)(x^2-x+1)(x^2+2x+4)"

4.

"16m^2-56mn+49n^2=(4m-7n)^2"

5.


"(x^2+1)^{-1\/2}(x-7)^2(x^2+1)^{1\/2}(x-7)"

"=(x-7)^3"